Circular No. 8347

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 H6 (SWAN) Preliminary parabolic orbital elements from the CCD astrometry on

IAUC 8346 (magnitudes from H = 7.5, n = 3):

T = 2004 May 12.716 TT  ω = 269.^{o}100
Ω = 317.219 }2000.0
q = 0.77619 AU  i = 107.657
2004TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
May 25 3^{h}28.^{m}40 −12^{o}53.^{′}6 1.391 0.812 35.^{o}2 46.^{o}0  7.5
30 3 19.57 −13 11.5 1.346 0.845 38.9 48.8  7.6
June 4 3 10.28 −13 21.6 1.291 0.887 43.2 51.5  7.7
9 3 00.35 −13 25.7 1.227 0.936 48.2 54.0  7.7
14 2 49.45 −13 25.1 1.156 0.991 53.8 55.9  7.8
19 2 37.12 −13 21.5 1.079 1.050 60.0 57.0  7.8
24 2 22.73 −13 15.6 0.998 1.112 67.0 57.3  7.8

SUPERNOVA 2004bx D. Magee, B. Holden, R. Bouwens, and G. Illingworth, University of

California, Santa Cruz; J. Blakeslee and H. Ford, Johns Hopkins University, and the Advanced Camera for Surveys (ACS) Science Team report the discovery of an apparent supernova in the galaxy cluster CL 1358+62, observed in exposures obtained with the Hubble Space Telescope’s ACS Wide Field Camera with the F850LP filter (at m_{AB} = 26.3 ± 0.2; ‘AB’ magnitudes are defined by Oke 1974, Ap.J. Suppl. 27, 21) on Apr. 6.82 UT, with the F775W filter (at m_{AB} = 25.67 ± 0.06) on Apr. 7.81, with the F850LP filter (at m_{AB} = 26.5 ± 0.3) on May 6.62, and with the F625W filter (at m_{AB} = 25.60 ± 0.05) on May 7.19. The new object is located at α = 13^{h}59^{m}51^{s}.06, δ = +62^{o}30′27′′.9 (equinox 2000.0), which is 0′′.6 east and 0′′.5 north of the center of the brightest nearby galaxy [having redshift 0.3174; cf. Fabricant et al. (1991, Ap.J. 831, 33)]. It is not clear if the nearby galaxy is the host galaxy or if the new object may reside in a background galaxy. SN 2004bx is not present in ACS observations from 2002 May 14 with the F775W filter (5σ limiting m_{AB} 27.9) and with the F850LP filter (5σ limiting m_{AB} 27.1).

2004 May 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green